OpenStudy (anonymous):

Identify the solution(s) of the system of equations, if any. 2x + 6y = 22 y = 5 - 3x

OpenStudy (anonymous):

equation 2 *-6 and add it to equation 1

OpenStudy (anonymous):

-6y= -30+18

OpenStudy (anonymous):

2x=22-30+18x

OpenStudy (anonymous):

x=.5

OpenStudy (anonymous):

and the other is 3.5 right?

OpenStudy (anonymous):

y=5-(3*.5)=3.5

OpenStudy (anonymous):

so it would read (3.5, 0.5)

OpenStudy (anonymous):

no....(0.5, 3.5) cuz we write it in xyz format

OpenStudy (anonymous):

(x,y)=(.5,3.5)

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

Ok....thanks!

OpenStudy (anonymous):

ur welcome